# Regurgitative Training: The Value of Real Data in Training Large Language Models

### v1

@@ 5.2 Mitigation Strategy based on Data Mixture
We first test this data mixture strategy under LLM fine-tuning setting. With three LLMs, there are 3 possible mixture configurations: (i) mixing GPT-3.5 with GPT-4, (ii) mixing GPT-3.5 with LLAMA2, and (iii) mixing GPT-4 with LLAMA2. We consider configurations (i) and (iii) in particular. Configuration (iii) mixes two top-of-the-line LLMs developed by different companies and, based on our explorations in Section 4, they exhibit different lexical diversity. In other words, this configuration is a more direct evaluation of the proposed strategy. However, it is unfortunately confounded by the fact that LLAMA2-generated data have higher translation error rates than GPT-generated data, so the performance outcomes of this configuration cannot be solely attributed to changes in lexical diversity levels. In comparison, configuration (i) is less confounded because the two GPT models have similar translation quality, although their mixture also bring in less additional lexical variations. Taken together, we present results from both configurations to offer a more comprehensive evaluation of the data mixture strategy.
@@
Interestingly, mixing data generated by GPT-4 and LLAMA2 results in performance levels that fall in between using the two constituent LLMs alone, but mixing data generated by GPT-3.5 and GPT-4 can match the performance of GPT-4 (the better-performing LLM of the two) and even slightly outperform it when twice the regurgitative training data are added. These results suggest that the effectiveness of the data mixture strategy is nuanced. When constituent LLMs differ both in terms of quality and lexical diversity, mixing their data together may not lead to better regurgitative training performance, because the potential benefit of greater lexical diversity is offset by having a more error-prone LLM in the mix. In the same vein, if two LLMs with similar quality yet different lexical diversity can be identified, their mixture can indeed mitigate performance loss of regurgitative training to some extent.
@@
We find highly consistent results as in the fine-tuning case. Mixing data generated by the two baseline models, which differ on quality, results in regurgitative training performance that falls between the two constituent models. Moreover, when training the high-performance baseline model with the data mixture (left plot of Figure 13), the performance follows a downward trend, likely because data generated by low-performance model substantially contaminate the quality of data mixture. Consequently, using twice the amount of mixture data in regurgitative training is worse in this case, as it further accelerates the performance decline. In contrast, mixing data generated by high-performance model and GPT-3.5, both of which have good quality, shows promising results. Regurgitative training with the data mixture matches the performance of training with real data when the same amount of data is used, and in fact outperforms it when twice the amount is used. The performance gain is especially evident when training the low-performance baseline model (right plot of Figure 14). Finally, we note that the data generated by high-performance model and GPT-3.5 have very similar quality (BLEU difference smaller than 0.3%). Therefore, the encouraging results from the high-performance model / GPT-3.5 mixture also lend additional support for our mechanism analyses in Section 4 – errors in generated data is not the only factor affecting regurgitative training performance, and other factors (such as lexical diversity) can be at play.
@@ 5.3 Mitigation Strategy based on AI Detection
Starting from the fine-tuning setting, we train an AI detection classifier using randomly sampled 75,000 German-English sentence pairs (not previously used in regurgitative fine-tuning) and their corresponding translations by GPT-3.5, GPT-4, and LLAMA2. For each LLM, we construct a balanced labeled dataset with 150,000 instances, where each of the 75,000 German sentences has exactly one real human-generated translation and one LLM-generated translation. Each data instance is structured as $input=(g_{1},g_{2},\ldots,g_{M},[SEP],e_{1},e_{2},\ldots,e_{N},[SEP]),label\in\{0,1\}$ where $label=0$ marks that the English translation is generated by an LLM and 1 otherwise. Same as how we have trained the BLEU prediction model in the previous section, we retrieve token embeddings from BERT. For each LLM, we train a separate classifier on 80% of labeled data and evaluate it on the remaining 20%. Performance scores of various supervised techniques are listed in Appendix C. The Linear Discriminant Analysis (LDA) turns out to have highest predictive performance for both GPT-3.5 and GPT-4, and Logistic Regression has the best performance for LAMMA2.
@@
We can see that regurgitative training in the order of resemblance with real data can indeed mitigate the performance loss for LLAMA2 and GPT-3.5, and the performance gain is larger for LLAMA2. This strategy, however, does not seem to be effective for GPT-4. We suspect such variations in mitigation effectiveness has to do with the capability of AI detection classifier – indeed, AI detection is most accurate for LLAMA2-generated data and least accurate for GPT-4-generated data (see Appendix C).
Next, we apply the same mitigation strategy on transformer models trained from scratch. Performance evaluations of various AI detection classifiers are again listed in Appendix C. We choose a Logistic Regression classifier for the high-performance baseline model and a LDA classifier for the low-performance baseline model. The regurgitative training results, with data ranked by class 1 predicted probabilities, are shown in Figure 16.
@@
These encouraging results highlight the promising utility of AI detection outside of its conventional use case. Besides identifying AI-generated content, a capable AI detector can also be re-purposed to guide more meaningful regurgitative training. Meanwhile, we acknowledge that this mitigation strategy clearly does not have unlimited capacity. Under both the fine-tuning setting and with a low-performance transformer model trained from scratch, even the AI detection-enhanced regurgitative training cannot catch up with the performance of training with real data.
@@ 6 Discussions
Our analyses in this paper give a negative answer to this question. Training a new LLM using data generated (at least partially) by other LLMs, a process we refer to as regurgitative training, generally results in lower performance than training with real data. While performance loss of regurgitative training has been documented in Shumailov et al., (2023) with early versions of generative models (non-transformer-based models or pre-trained models before GPT-3.5), our work provides more comprehensive evidence by both fine-tuning commercial LLMs (including GPT-3.5, GPT-4, and LLAMA2) and training small-scale transformer models from scratch. Our explorations reveal more nuanced performance effects of regurgitative training. Under both fine-tuning and training-from-scratch settings, regurgitative training with data generated by a competent model may still improve performance to a small extent and at a much lower speed / magnitude than training with real data, whereas regurgitative training with data generated by a poor model hurts performance. These effects manifest even when only a small proportion of training data are synthetic. Even in the rare case where regurgitative training outperforms training with real data (i.e., training low-performance model with high-performance model generated data, Figure 4), such advantage disappears after a large amount of synthetic data is used.
@@
These explorations of underlying mechanisms also produce potential strategies to mitigate performance loss of regurgitative training. In total, we propose and test three different strategies, respectively designed to address the issues of data quality / error, lack of lexical diversity, and both. The quality-based mitigation strategy prioritizes the use of high-quality data for regurgitative training, where “quality” can be gauged either by prediction confidence or via a supervised learning approach. The data mixture strategy seeks to enhance lexical diversity by mixing together data generated from different LLMs. The AI detection-based strategy re-purposes an AI detection classifier to identify LLM-generate data that resemble real data, then prioritize their use in regurgitative training. While all three strategies can reduce performance loss to some extent, their relative effectiveness also demonstrates some interesting nuances. First, they tend to be more effective on transformer models trained from scratch than fine-tuned LLMs. Under the fine-tuning setting, none of the migration strategies can bridge the performance gap between regurgitative training and training with real data; in contrast, data mixture and AI detection based strategies can outperform training with real data on models trained from scratch. Second, the data mixture strategy does not perform well if constituent LLMs differ both in terms of quality and lexical diversity – the drop in data quality offsets the benefits of increased diversity. Instead, this strategy performs much better if constituent LLMs have comparable quality but still contribute diversity benefits (e.g., using competent models trained on different data or architectures). Finally, success of the AI detection strategy hinges on the ability to differentiate LLM- vs. human-generated data. Greater capable in AI detection generally results in better regurgitative training performance.
@@
Several implications for both researchers and practitioners working with LLMs are worth noting. First, we urge caution when utilizing synthetic LLM-generated data when training or fine-tuning LLMs. Despite the multitude of amazing capabilities of LLMs, at their current stage, regurgitative training cannot create sustained performance improvement. Therefore, datasets that are organically generated and carefully curated (such as the Europarl corpus for machine translation and SQuAD corpus for Q&A) remain part of the core assets of LLM development. Second, the prevalent use of LLMs implies that LLM-generated data would likely take up a non-trivial proportion of online content in the near future, and some degree of regurgitative training may be unavoidable. Recognizing this trend, we advocate for more careful use of LLM-generated data. Our results suggest that “data quality” trumps “data quantity” in regurgitative training – it is generally more advantageous to use data with higher prediction confidence, greater linguistic richness, and higher resemblance to real data than merely using a larger quantity of data with questionable quality. Finally, the baseline performance of the model being regurgitatively trained also matters. All else being equal, a more capable baseline model suffers less performance loss due to regurgitative training. Therefore, it is important for researchers and businesses to first thoroughly train their baseline models before engaging in regurgitative training, which helps control the adverse impact of regurgitative training.
@@
Our work also opens up a few interesting future research directions. Capabilities and performance of modern LLMs are constantly evolving. Is the negative performance impact of regurgitative training just a transient pattern reflecting limitations of available LLMs (which may disappear as more powerful LLMs are created in the future), or is it a fundamental issue of current paradigm of generative AI? Existing work such as Shumailov et al., (2023) and Gerstgrasser et al., (2024) attempt to answer this question by resorting to analyzing simplified models (e.g., one-dimensional Gaussian processes). Future work can employ more advanced theoretical frameworks to derive deeper understandings. Moreover, effective regurgitative training with synthetic data represents an emerging field of increasing importance. Our proposed mitigation strategies are only the first steps rather than final words, and we encourage future work to design more potent strategies that can be adopted in practice. We believe more productive use of synthetic data for LLM training requires both theoretical understanding of its performance upper bound, as well as practical algorithms and methods to achieve what is possible.
Appendix A Identifying Low-/High-Performance Models in Q&A Task
@@
Appendix B BLEU Prediction Performance in Quality-Based Mitigation Strategy
@@
Model MAE MSE RMSE
Bayesian Ridge 0.1602 0.0415 0.2038
Ridge Regression 0.1601 0.0416 0.2038
Linear Regression 0.1602 0.0416 0.2039
Light Gradient Boosting Machine 0.1608 0.0417 0.2043
Orthogonal Matching Pursuit 0.1632 0.0428 0.2069
Extra Trees Regressor 0.1613 0.0421 0.2051
Gradient Boosting Regressor 0.1636 0.0427 0.2067
Extreme Gradient Boosting 0.1617 0.0430 0.2074
@@
Model MAE MSE RMSE
Bayesian Ridge 0.1627 0.0429 0.2072
Ridge Regression 0.1627 0.0430 0.2072
Linear Regression 0.1627 0.0430 0.2073
Light Gradient Boosting Machine 0.1631 0.0430 0.2073
Orthogonal Matching Pursuit 0.1657 0.0442 0.2103
Extra Trees Regressor 0.1636 0.0434 0.2083
Gradient Boosting Regressor 0.1659 0.0440 0.2098
Extreme Gradient Boosting 0.1643 0.0445 0.2109
@@
Model MAE MSE RMSE
Bayesian Ridge 0.1350 0.0301 0.1735
Ridge Regression 0.1349 0.0301 0.1735
Linear Regression 0.1349 0.0301 0.1736
Light Gradient Boosting Machine 0.1374 0.0306 0.1749
Orthogonal Matching Pursuit 0.1381 0.0311 0.1764
Extra Trees Regressor 0.1397 0.0314 0.1771
Gradient Boosting Regressor 0.1400 0.0314 0.1772
Extreme Gradient Boosting 0.1371 0.0314 0.1772
Appendix C Performance Evaluation of AI Detection Classifiers

### v2

@@ 5.2 Mitigation Strategy based on Data Mixture
We first test this data mixture strategy under LLM fine-tuning setting. With three LLMs, there are 3 possible mixture configurations: (i) mixing GPT-3.5 with GPT-4, (ii) mixing GPT-3.5 with LLAMA2, and (iii) mixing GPT-4 with LLAMA2. We consider configurations (i) and (iii) in particular. Configuration (iii) mixes two top-of-the-line LLMs developed by different companies and, based on our explorations in Section 4, they exhibit different lexical diversity. In other words, this configuration is a more direct evaluation of the proposed strategy. However, as we have shown before, LLAMA2-generated data have higher translation error rates than GPT-generated data, so the performance outcomes of this configuration may not be solely attributed to changes in lexical diversity levels. In comparison, configuration (i) is less confounded because the two GPT models have similar translation quality, although their mixture also bring in less additional lexical variations. Taken together, we present results from both configurations to offer a more comprehensive evaluation of the data mixture strategy.
@@
Interestingly, mixing data generated by GPT-4 and LLAMA2 results in performance levels that fall in between using the two constituent LLMs alone, but mixing data generated by GPT-3.5 and GPT-4 can match the performance of GPT-4 (the better-performing LLM of the two) and even slightly outperform it when twice the regurgitative training data are added. These results suggest that the effectiveness of the data mixture strategy is nuanced. When constituent LLMs differ both in terms of quality and lexical diversity, mixing their data together may not lead to better regurgitative training performance, because the potential benefit of greater lexical diversity is offset by having a more error-prone LLM in the mix. In the same vein, if two LLMs with similar quality yet different levels of lexical diversity can be identified, their mixture can indeed mitigate performance loss of regurgitative training to some extent.
@@
We find highly consistent results as in the fine-tuning case. Mixing data generated by the two baseline models, which differ on quality, results in regurgitative training performance that falls between the two constituent models. Moreover, when training the high-performance baseline model with the data mixture (left plot of Figure 13), the performance follows a downward trend, likely because data generated by low-performance model substantially contaminate the quality of data mixture. Consequently, using twice the amount of mixture data in regurgitative training is worse in this case, as it further accelerates the performance decline. In contrast, mixing data generated by high-performance model and GPT-3.5, both of which have good quality, results in higher performance than both constituent models. In fact, regurgitative training with the data mixture matches the performance of training with real data when the same amount of data is used, and in fact outperforms it when twice the amount is used. The performance gain is especially evident when training the low-performance baseline model (right plot of Figure 14). Finally, we note that the data generated by high-performance model and GPT-3.5 have very similar quality (BLEU difference smaller than 0.3%). Therefore, the encouraging results from the high-performance model / GPT-3.5 mixture also lend additional support for our mechanism analyses in Section 4 – errors in generated data are not the only factor affecting regurgitative training performance, and other factors (such as lexical diversity) can be at play.
@@ 5.3 Mitigation Strategy based on AI Detection
Starting from the fine-tuning setting, we train an AI detection classifier using randomly sampled 75,000 real translations and another 75,000 translations respectively generated by GPT-3.5, GPT-4, and LLAMA2 (not previously used in regurgitative fine-tuning). For each LLM, we construct a balanced labeled dataset with 150,000 instances, half of which are real human-generated translations and the other half are LLM-generated translations. Each data instance is structured as $input=(g_{1},g_{2},\ldots,g_{M},[SEP],e_{1},e_{2},\ldots,e_{N},[SEP]),label\in\{0,1\}$ where $label=0$ marks that the English translation is generated by an LLM and 1 otherwise. Same as how we have trained the BLEU prediction model in the previous section, we retrieve token embeddings from multilingual BERT. For each LLM, we train a separate classifier on 80% of labeled data and evaluate it on the remaining 20%. Performance scores of various supervised techniques are listed in Appendix E. The Linear Discriminant Analysis (LDA) turns out to have highest predictive performance for both GPT-3.5 and GPT-4, and Logistic Regression has the best performance for LAMMA2.
@@
We can see that regurgitative training in the order of resemblance with real data can indeed mitigate the performance loss for LLAMA2 and GPT-3.5, and the performance gain is larger for LLAMA2. This strategy, however, does not seem to be effective for GPT-4. We suspect such variations in mitigation effectiveness have to do with the capability of AI detection classifier – indeed, AI detection is most accurate for LLAMA2-generated data and least accurate for GPT-4-generated data (see Appendix E).
Next, we apply the same mitigation strategy on transformer models trained from scratch. Performance evaluations of various AI detection classifiers are again listed in Appendix E. We choose a Logistic Regression classifier for the low-performance baseline model and a LDA classifier for the high-performance baseline model. The regurgitative training results, with data ranked by class 1 predicted probabilities, are shown in Figure 16.
@@
These encouraging results highlight the promising utility of AI detection outside of its conventional use case. Besides identifying AI-generated content, a capable AI detector can also be re-purposed to guide more meaningful regurgitative training. Meanwhile, we acknowledge that this mitigation strategy clearly does not have unlimited capacity. If synthetic data are generated by a less capable model (specifically, LLAMA2 in the fine-tuning setting and the low-performance model in the training-from-scratch setting), even the AI detection-enhanced regurgitative training (marked by purple lines in Figures 15-16) cannot catch up with the performance of training with real data.
@@ 6 Discussions
Our analyses in this paper give a negative answer to this question. Training a new LLM using data generated (at least partially) by itself or other LLMs, a process we refer to as regurgitative training, generally results in lower performance than training with real data. While performance loss of regurgitative training has been documented in Shumailov et al., (2023) with early versions of generative models (non-transformer-based models or pre-trained models before GPT-3.5), our work provides more comprehensive evidence by both fine-tuning commercial LLMs (including GPT-3.5, GPT-4, and LLAMA2) and training small-scale transformer models from scratch. Our explorations also reveal more nuanced performance effects of regurgitative training. Under both fine-tuning and training-from-scratch settings, regurgitative training with data generated by a competent model may still improve performance to a small extent (compared to the baseline performance without regurgitative training), but such performance improvement typically comes in a much lower speed / magnitude than training with real data. In contrast, regurgitative training with data generated by a poor model hurts performance (and of course also underperforms training with real data). These effects manifest even when only a small proportion of training data are synthetic. Even in the rare case where regurgitative training outperforms training with real data (i.e., training low-performance model with high-performance model generated data, Figure 4), such advantage disappears after a large amount of synthetic data is used.
@@
These explorations of underlying mechanisms also produce potential strategies to mitigate performance loss of regurgitative training. In total, we propose and test three different strategies, respectively designed to address the issues of data quality / error, lack of lexical diversity, and both. The quality-based mitigation strategy prioritizes the use of high-quality data for regurgitative training, where “quality” can be gauged either by prediction confidence or via a supervised learning approach. The data mixture strategy seeks to enhance lexical diversity by mixing together data generated from different LLMs. The AI detection-based strategy re-purposes an AI detection classifier to identify LLM-generated data that resemble real data, then prioritize their use in regurgitative training. While all three strategies can reduce performance loss to some extent, their relative effectiveness demonstrates some interesting nuances. First, they tend to be more effective on transformer models trained from scratch than fine-tuned LLMs. Under the fine-tuning setting, none of the migration strategies can bridge the performance gap between regurgitative training and training with real data; in contrast, when applied on the low-performance baseline model, the data mixture and AI detection based strategies can outperform training with real data on models trained from scratch. Second, the data mixture strategy does not perform well if constituent LLMs differ both in terms of quality and lexical diversity – the drop in data quality offsets the benefits of increased diversity. Instead, this strategy performs much better if constituent LLMs have comparable quality but still contribute diversity benefits (e.g., using competent models trained on different data or architectures). Finally, success of the AI detection strategy hinges on the ability to differentiate LLM- vs. human-generated data. Greater capability in AI detection generally results in better regurgitative training performance.
@@
Several implications for both researchers and practitioners working with LLMs are worth noting. First, we urge caution when utilizing synthetic LLM-generated data when training or fine-tuning LLMs. Despite the multitude of amazing capabilities of LLMs, at their current stage, regurgitative training cannot create sustained performance improvement and often deteriorates the model’s performance. Therefore, datasets that are organically generated and carefully curated (such as the Europarl corpus for machine translation and SQuAD corpus for Q&A) remain part of the core assets of LLM development. Second, the prevalent use of LLMs implies that LLM-generated data would likely take up a non-trivial proportion of online content in the near future, and some degree of regurgitative training may be unavoidable. Recognizing this trend, we advocate for more careful use of LLM-generated data. Our results suggest that “data quality” trumps “data quantity” in regurgitative training – it is generally more advantageous to use data with higher prediction confidence, greater linguistic richness, and higher resemblance to real data than merely using a larger quantity of data with questionable quality. Moreover, the baseline performance of the model being regurgitatively trained also matters. All else being equal, a more capable baseline model suffers less performance loss due to regurgitative training. Therefore, it is important for researchers and businesses to first thoroughly train their baseline models before engaging in regurgitative training, which helps control the adverse impact of regurgitative training. Finally, human-generated data often reflect human needs, preferences, and value judgments (e.g., online reviews reflect user preferences about products), whereas LLM-generated data may not necessarily reflect the same degree of genuine human needs and value judgments. Consequently, regurgitative training can also impact a business’ ability to transform data into value. Assessing the nature of this impact and designing mitigation strategies to manage it is also an important organizational objective.
@@
Our work also opens up a few interesting future research directions. Capabilities and performance of modern LLMs are constantly evolving. Is the negative performance impact of regurgitative training just a transient pattern reflecting limitations of available LLMs (which may disappear as more powerful LLMs are created in the future), or is it a fundamental issue of current paradigm of generative AI? Existing work such as Shumailov et al., (2023) and Gerstgrasser et al., (2024) attempt to answer this question by resorting to analyzing simplified models (e.g., one-dimensional Gaussian processes). Future work can employ more advanced theoretical frameworks to derive deeper understandings. Moreover, effective regurgitative training with synthetic data represents an emerging field of increasing importance. Our proposed mitigation strategies are only the first steps rather than final words, and we encourage future work to design more potent strategies that can be adopted in practice. We believe more productive use of synthetic data for LLM training requires both theoretical understanding of its performance upper bound, as well as practical algorithms and methods to achieve what is possible. Lastly, LLMs are used not only in tasks with established performance metrics (such as the machine translation and Q&A tasks considered in this work), but also in open-ended tasks whose quality are very hard to gauge (e.g., creative ideation). How to conceptually think about and empirically evaluate the impact of regurgitative training in these open-ended tasks remains an interesting question.
Appendix A Robustness Check: Regurgitative Training with A Fixed Percentage of Training Data
Figure 17 shows the results of regurgitative training on both low- and high-performance baseline models with model-generated data that amount to 10% the size of their respective training data. Figure 18 shows the results when a certain percentage of real data is mixed in.
Figure 17: Performance of Regurgitative Training Transformer Models with A Fixed Percentage of Training Data (two plots have different $y$-axis scales for better readability)
Figure 18: Performance of Regurgitative Training Transformer Models with Different Proportions of Real Data and A Fixed Percentage of Training Data (two plots have different $y$-axis scales for better readability)
Appendix B Identifying Low-/High-Performance Models in Q&A Task
@@
Appendix C BLEU Prediction Performance in Quality-Based Mitigation Strategy
@@
Model MSE RMSE MAE
Bayesian Ridge 0.0415 0.2038 0.1602
Ridge Regression 0.0416 0.2038 0.1601
Linear Regression 0.0416 0.2039 0.1602
Light Gradient Boosting Machine 0.0417 0.2043 0.1608
Orthogonal Matching Pursuit 0.0428 0.2069 0.1632
Extra Trees Regressor 0.0421 0.2051 0.1613
Gradient Boosting Regressor 0.0427 0.2067 0.1636
Extreme Gradient Boosting 0.0430 0.2074 0.1617
@@
Model MSE RMSE MAE
Bayesian Ridge 0.0429 0.2072 0.1627
Ridge Regression 0.0430 0.2072 0.1627
Linear Regression 0.0430 0.2073 0.1627
Light Gradient Boosting Machine 0.0430 0.2073 0.1631
Orthogonal Matching Pursuit 0.0442 0.2103 0.1657
Extra Trees Regressor 0.0434 0.2083 0.1636
Gradient Boosting Regressor 0.0440 0.2098 0.1659
Extreme Gradient Boosting 0.0445 0.2109 0.1643
@@
Model MSE RMSE MAE
Bayesian Ridge 0.0301 0.1735 0.1350
Ridge Regression 0.0301 0.1735 0.1350
Linear Regression 0.0301 0.1736 0.1349
Light Gradient Boosting Machine 0.0306 0.1749 0.1374
Orthogonal Matching Pursuit 0.0311 0.1764 0.1381
Extra Trees Regressor 0.0314 0.1771 0.1397
Gradient Boosting Regressor 0.0314 0.1772 0.1400
Extreme Gradient Boosting 0.0314 0.1772 0.1371
Appendix D Quality-Based Mitigation Strategy on Q&A Task: Ranking by Average Entropy
In the following Figure 20, we report the evaluation results of quality-based mitigation strategy on the Q&A Task, where we prioritize articles based on the average entropy value of their constituent answers (rather than the lowest entropy value).
Figure 20: Quality-Based Mitigation Strategy: Results on Q&A Task (Ranking by Average Entropy)
Appendix E Performance Evaluation of AI Detection Classifiers